The Dahlia Company has net income of $155,850. There are currently 27.85 days sales in receivables. Total assets are $844,000, total receivables are $146,900, and the debt-equity ratio is .45 What is the companys profit margin? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Profit margin What is the companys total asset turnover? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Total asset turnover What is the companys ROE? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) ROE times
Answer #1

days sales in receivables = [ Receivables / Sales ] * 365 days

27.85 days = [ $146,900 / Sales ] * 365 days

Sales = $146,900 * 365 days / 27.85 days = $1,925,260

Profit margin = Net income / Sales = $155,850 / 1,925,260 = 8.10%

Total asset turnover = Sales / Total assets = $1,925,260 / $844,000 = 2.28 times

ROE = Profit margin * Total asset turnover * ( 1 + Debt equity ratio ) = 8.10% * 2.28 * ( 1 + 0.45 ) = 26.78%